double integral of xdydx +2ydzdx+ 3zdxdy over the surface of the spherex^2 +y^2 +z^2 is



1
Expert's answer
2021-02-25T01:32:27-0500

Given integral is-

"\int \int xdxdy+\int \int 2ydzdx+\int \int3zdxdy"


Given Region is "x^2+y^2+z^2=1"


Let's evaluate the first term- "\int\int xdxdy"

Putting "z=0, \text { we get } x^2+y^2=1\implies y=\sqrt{1-x^2}"


"=\int_0^1\int_0^{\sqrt{1-x^2}} xdxdy"


"=\int_0^1x[y]_0^{\sqrt{1-x^2}}dx"


"=\int_0^1x(\sqrt{1-x^2})dx"



"\text{ Let }1-x^2=t"

"\\ -2xdx=dt\\ xdx=\dfrac{dt}{-2}"

"=\int_1^0\dfrac{\sqrt{t}}{-2}dt"


"=-\dfrac{1}{2}[\dfrac{t^\frac{3}{2}}{\frac{3}{2}}]_1^0"


"=\dfrac{1}{3}(1)=\dfrac{1}{3}~~~~~~~-(1)"


Solving integral of second term-


"\int\int_R 2ydxdz"


Putting the value of "y=\sqrt{1-x^2-z^2}" in above integral as


"=2\int_0^1\int_0^{\sqrt{1-z^2}}\sqrt{1-z^2-x^2}dzdx"


"=2\int_0^1(\dfrac{x}{2}\sqrt{1-z^2-x^2}+\dfrac{1-z^2}{2}sin^{-1}(\dfrac{x}{\sqrt{1-z^2}})_0^{\sqrt{1-z^2}})dz"


"=2\int_0^1(\dfrac{1-z^2}{2}sin^{-1}(\dfrac{\sqrt{1-z^2}}{\sqrt{1-z^2}})-0)dz"


"=2\int_0^1\dfrac{1-z^2}{2}sin^{-1}(1)dz"


"=\int_0^1(1-z^2)\dfrac{\pi}{2}dz"


"=\dfrac{\pi}{2}(z-\dfrac{z^3}{3})_0^1"


"=\dfrac{\pi}{2}(1-\dfrac{1}{3})=\dfrac{\pi}{2}(\dfrac{2}{3})=\dfrac{\pi}{3}~~~~~~~~~-(2)"



Now integrate the third term-

"\int\int_R3zdxdy"


Now Putting the value "z =\sqrt{1-x^2-y^2}" and applying the same procedure as above-


"=3\int_0^1\int_0^{\sqrt{1-y^2}}\sqrt{1-y^2-x^2}dxdy"


"=3\int_0^1(\dfrac{x}{2}\sqrt{1-y^2-x^2}+\dfrac{1-y^2}{2}sin^{-1}(\dfrac{x}{\sqrt{1-y^2}})_0^{\sqrt{1-y^2}})dy"


"=3\int_0^1(\dfrac{1-y^2}{2}sin^{-1}(\dfrac{\sqrt{1-y^2}}{\sqrt{1-y^2}})-0)dy"


"=3\int_0^1\dfrac{1-y^2}{2}sin^{-1}(1)dy"


"=\dfrac{3}{2}\int_0^1(1-y^2)\dfrac{\pi}{2}dy"


"=\dfrac{3\pi}{4}(y-\dfrac{y^3}{3})_0^1"


"=\dfrac{3\pi}{4}(1-\dfrac{1}{3})=\dfrac{3\pi}{4}(\dfrac{2}{3})=\dfrac{\pi}{2}~~~~~~~~~-(3)"



Adding the three equation we get-


"\int \int xdxdy+\int \int 2ydzdx+\int \int3zdxdy=\dfrac{1}{3}+\dfrac{\pi}{3}+\dfrac{\pi}{2}=\dfrac{5\pi}{6}+\dfrac{1}{3}"